OpenStudy (anonymous):

Solve 4(x + 1) + 8 = 24. 3.75 −2 5 3

OpenStudy (texaschic101):

4(x + 1) + 8 = 24 --- distribute through the parenthesis 4x + 4 + 8 = 24 -- simplify 4x + 12 = 24 --- subtract 12 from both sides 4x = 24 - 12 4x = 12 -- divide both sides by 4 x = 3 check.. 4(x + 1) + 8 = 24 4(3 + 1) + 8 = 24 4(4) + 8 = 24 16 + 8 = 24 24 = 24 (correct) so x = 3

OpenStudy (anonymous):

Jeez. how do you do that so fast?

OpenStudy (anonymous):

Opening the parenthesis is not necessary \[4(x+1)+8=24\] Subtract 8 \[4(x+1)+8-8=24-8\]\[4(x+1)=16\] Divide by 4 \[\frac{4(x+1)}{4}=\frac{16}{4}\]\[x+1=4\] Subtract 1 \[x+1-1=4-1\]\[x=3\]